OpenStudy (jacksonjrb):

Need help with a radical equation. Open question for equation. Will give medals.

OpenStudy (jacksonjrb):

\[\sqrt{2x+9}-\sqrt{x+1}=\sqrt{x+4}\]

OpenStudy (mokeira):

step 1, square to remove square root 

OpenStudy (anonymous):

for the left hand side, it should be in the form...\[(\sqrt{a}-\sqrt{b})^2=a-2\sqrt{ab}+b\]